40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 15319     Entry time: Wed May 6 00:31:09 2020
Author: gautam 
Type: Update 
Category: ALS 
Subject: Optomechanics during CARM offset reduction 

Summary:

The apparent increase in the ALS noise (witnessed in-loop, e.g. Attachment #2 here) during the CARM offset reduction may have an optomechanical origin. 

Details:

  • A simplified CARM plant was setup in Finesse - 3 mirror coupled cavity with PRM, ITM and ETM, 40m params for R/T/L used. 
  • For a sanity check, DC power buildup and coincident resonance of the PRC and arm cavity were checked. PRG and CARM linewidth also checks out, and scales as expected with arm losses.
  • To investigate possible optomechanical issues - I cut the input power to 300 mW (I estimate 600 mW incident on the PRM), set a PRG of ~20, to mimic what we have right now.
  • I drive the ITM at various CARM offsets, and measure the m/m transfer function to itself and the ETM.
  • Attachement #1 shows the results. 

Interpretation:

  • ericq had similar plots in his thesis, but I don't think the full implications of this effect were investigated, the context there was different.
  • The optomechanical resonance builds up at ~10 Hz and sweeps up to ~100 Hz as the CARM offset approaches zero, with amplification close to x100 at the resonance.
  • What this means is that the arm cavity is moving by up to 100x the ambient seismically driven dispalcements. 
  • The EX/EY uPDH servos have considerable gain at these frequencies, and so the AUX laser frequency can keep up with this increased motion (to be quantified exactly what the increase in residual is).
  • However, the ALS loop that maintains the frequency offset b/w the PSL and the AUX lasers is digitial, and only has ~20 dB gain at 30 Hz. - so the error signal for CARM control becomes noisier as we see.
  • I speculate that the multiple peaky features in the in-loop error signal are a result of some dynamical effects which Finesse presumably does not simulate.
  • The other puzzler is: this simulation would suggest that approaching the zero CARM offset from the other side (anti-spring) wouldn't have such instabilities building up. However, I am reasonably sure I've seen this effect approaching zero from both sides, though I haven't checked in the last month.
  • Anyways, if this hypothesis is correct, we can't really take advantage of the ~8pm RMS residual noise performance of the IR ALS system sadly, because of our 250g mirrors and 800mW input power
  • Possible workarounds:
    • High BW ALS - this would give us more gain at ~30 Hz and this wouldn't be a problem anymore really. But in my trials, I think I found that the IN2 gain on the CM board has to be inverted for this to work (the IN1 path and the IN2 path share a common AO path polarity, and we need the two paths to have the opposite polarity).
    • Cut the input power - this would reduce the optomechanical action, but presumably the vertex locking becomes noisier. In any case, this isn't really practical without some kind of motorized/remote-controlled waveplate for power adjustment. 

Update 415pm 5/6: Per the discussion at the meeting, I have now uploaded as Attachment #2 the force-->displacement (i.e. m/N) transfer functions. I now think these are appropriate units. For the ALS case, we could convert the m/N to Hz/N of extra frequency noise imprinted on the AUX laser due to the increased cavity motion. Is W/N really better here, since the mechanism is extra frequency noise on a beatnote, and there isn't really a PDH or DC error signal?

Attachment 1: CARMplant.pdf  207 kB  Uploaded Wed May 6 01:47:27 2020  | Hide | Hide all
CARMplant.pdf
Attachment 2: CARMplant_force2disp.pdf  280 kB  Uploaded Wed May 6 17:21:53 2020  | Hide | Hide all
CARMplant_force2disp.pdf
ELOG V3.1.3-